OpenStudy (anonymous):

Find the area and the circumference of a circle with diameter 6 ft. Use the value 3.14 for pi, and do not round your answers.

jimthompson5910 (jim_thompson5910):

You'll use these formulas A = pi*r^2 C = pi*d A = area C = circumference r = radius (half of the diameter) d = diameter pi = 3.14 approximately

jimthompson5910 (jim_thompson5910):

keep in mind that r^2 is not the same as r*2

OpenStudy (anonymous):

So for the circumference, 3.14*6?

OpenStudy (anonymous):

Is the area 9ft^2 ?

jimthompson5910 (jim_thompson5910):

area would be 9*pi = 9*3.14 = ??

OpenStudy (anonymous):

28.26?

jimthompson5910 (jim_thompson5910):

yep that's the area

OpenStudy (anonymous):

18.84 is the circumference?

OpenStudy (mathmale):

@Snufflebug02: Please be sure to include the correct units of measurement with your answers. 18.84 should not be left standing alone. Neither should 28.26.

OpenStudy (anonymous):

18.84 ft^2 and 28.26 ft^2

OpenStudy (anonymous):

Is that correct?

jimthompson5910 (jim_thompson5910):

`18.84 ft^2` is partially correct the 18.84 is correct, but the units for circumference are going to be ft instead of ft^2

OpenStudy (anonymous):

Okay thank you! I have another question, if you have time :)

jimthompson5910 (jim_thompson5910):

imagine you cut the circle and stretched it out to form a line like you see in this animation https://upload.wikimedia.org/wikipedia/commons/6/67/2pi-unrolled.gif

OpenStudy (anonymous):

A rectangle is placed around a semicircle as shown below. The width of the rectangle is 9 cm. Find the area of the shaded region. Use the value 3.14 for pi, and do not round your answer. Be sure to include the correct unit in your answer.

OpenStudy (anonymous):

Yes I understand now. Thank you for talking me through it

jimthompson5910 (jim_thompson5910):

`A rectangle is placed around a semicircle as shown below` you forgot to post the image

OpenStudy (anonymous):

I tried. I'm not sure how too

jimthompson5910 (jim_thompson5910):

click the "attach file" button or host the image on a site like imgur then post the link

OpenStudy (anonymous):

I don't see an "attach file" button. how do i use imgur?

jimthompson5910 (jim_thompson5910):

you go to http://imgur.com/ then drag and drop your image to upload it then share the link here

OpenStudy (anonymous):

It's not letting me. I'll try to explain it. Its a rectangle with a semicircle inside of it. And the top right and left corners are shaded in. That is the area that we need to find

OpenStudy (anonymous):

I'm sorry if this is difficult

jimthompson5910 (jim_thompson5910):

like this?